An electron acquires 6.90×10-16 J of kinetic energy when it is accelerated by an electric field from plate A to plate B.

What is the potential difference between the plates?

\Delta V_{\rm BA} = _________________ V

Which plate is at the higher potential?
Answer #1
Concepts and reason

This problem is based on the concept of work done by an electric field on a charge and work-energy theorem.

The work done is to be estimated by the provided kinetic energy. Then it is to be used to calculate the potential. The plate having higher potential is to be determined.

Fundamentals

Work energy Theorem: The change in the kinetic energy is equal to work done.

Write the expression for the work done on a charge in an electric field.

W=qVW = qV

Here, VV is potential, qq is charge and WW is work done.

Write the expression for the work-energy theorem.

W=ΔKW = \Delta K

Here, ΔK\Delta K is change in kinetic energy.

Write the expression for the work-energy theorem.

W=ΔKW = \Delta K

Substitute 6.90×1016J6.90 \times {10^{ - 16}}\;{\rm{J}} for ΔK\Delta K.

W=6.90×1016JW = 6.90 \times {10^{ - 16}}\;{\rm{J}}

Write the expression for the work done on a charge in an electric field.

W=qVW = qV

V=WqV = \frac{W}{q}

Substitute 6.90×1016J6.90 \times {10^{ - 16}}\;{\rm{J}} for WW and 1.6×1019C1.6 \times {10^{ - 19}}\;{\rm{C}} for qq.

V=6.90×1016J1.6×1019C=4313V\begin{array}{c}\\V = \frac{{6.90 \times {{10}^{ - 16}}\;{\rm{J}}}}{{1.6 \times {{10}^{ - 19}}\;{\rm{C}}}}\\\\ = 4313{\rm{ V}}\\\end{array}

The electron has negative charge and it is accelerated from plate AA to plate BB.Therefore the plate BBwill have higher potential.


Ans:

The potential between the plates is 4313V4313{\rm{ V}} and plate BB has higher potential.